OpenStudy (f_jayyy):

What is the remainder when x 3 - 7x - 6 is divided by the linear factor x - 4?

OpenStudy (tkhunny):

Have you considered Synthetic Division?

OpenStudy (f_jayyy):

@Luigi0210 @Jaynator495 @AloneS @Awolflover1 @.Sam. @narissa

OpenStudy (f_jayyy):

can you explain what that is to me? @tkhunny

OpenStudy (tkhunny):

Don't tag the world. Show your work, instead.

OpenStudy (tkhunny):

You've never seen Synthetic Division?

OpenStudy (f_jayyy):

no i havent

OpenStudy (karleighxxvictoria):

i got -4x-6 over x-4

OpenStudy (f_jayyy):

but we are looking for the remainder @karleighxxvictoria

OpenStudy (steve816):

If you want, long division is an option.

OpenStudy (karleighxxvictoria):

ohhhh I see nevermind i was wrong!

OpenStudy (tkhunny):

1) Start by writing down the coefficients of ALL the powers of x. Include those that are zero. 1 0 -7 -6 in this case. 2) Pull down the first number and just keep multiplying by the value from the factor. We have "x-4", so use 4. 

OpenStudy (tkhunny):

The values in the bottom come from the SUM of the two values directly above. The values in the middle come from multiplying each value from the bottom, in order, by the value on the left. Anyway, that's one way to do it.

OpenStudy (f_jayyy):

oh okay cool thx

OpenStudy (tkhunny):

But what does that tell us?

OpenStudy (tkhunny):

Again, just read the coefficients \(\dfrac{x^{3} - 7x - 6}{x-4} = x^{2} + 4x + 9 + \dfrac{30}{x-4}\) What is the "Remainder" that you seek?